What is y in this equation 3/2y - y =4 + 1/2?

Answer by Alan3354(69443)

3/2y - y =4 + 1/2
I'll assume you mean (3/2)y for the 1st term.
y/2 = 9/2
y = 9
----------
It could be 3/(2y), which would make a difference. Parentheses are free, use some.